OpenStudy (anonymous):

what is the square rout of 4/25, 1/4, 1/121

OpenStudy (angela210793):

sqr(4/25)=+-2/5 sqr(1/4)=+-1/2 sqr(1/121)=+-1/11

OpenStudy (anonymous):

\[\sqrt{\frac{4}{25}}=\frac{\sqrt{4}}{\sqrt{25}}= \frac{2}{5}\]

OpenStudy (anonymous):

+-

OpenStudy (anonymous):

although does it need a -

OpenStudy (anonymous):

not sure on the rules of square root

OpenStudy (anonymous):

thanks

OpenStudy (angela210793):



OpenStudy (anonymous):

hmm yeah i know but there was a proof yesterday that kushawasha posted with imaginary.. that is probably why

OpenStudy (angela210793):

oh...i saw tht...was pretty interesting :)